Please help me. I have no idea what this means.
Approximate square root 31 by giving the two whole numbers that it lies between
Thanks so much for your help.

Answer by edjones(7569)

Approximate square root 31
sqrt(31) is between 5 & 6
Guess: 5.5
Divide: 31/5.5=5.6
Average:(5.5+5.6)/2=5.55
Divide: 31/5.55=5.586
average: (5.55+5.586)/2=5.705
divide: 31/5.705=5.4329
average: (5.4329+5.705)/2=5.5675
divide:31/5.5675=5.56803
average: (5.5675+5.56803)/2=5.56777
.
You can go on indefinately dividing and averaging, getting a closer approximation of the answer. Since all square roots that are not integers are irrational numbers we will never find an exact answer.
.
My calculator finds: 5.56776... so we did pretty well.
.
Ed